# A Survey on Data Selection for LLM Instruction Tuning

### v1

@@ 3.4 Methods based on Small Models
4 Evaluation methods and Result Analysis
The effectiveness of data selection methods depends on the quality of the subset filtered from the given dataset. To measure the quality of the subset, the LLMs fine-tuned on the subset are evaluated on different benchmarks in different approaches, which can be divided into three categories: wining rate, inner comparison and external comparison.
4.1 Wining Rate
To evaluate the effectiveness of the dataset selection methods, the winning rate of the LLM-sub compared to the base LLM is calculated as:
$(Num(win)-Num(lose))/Num(all)+1$ (10)
LLM-sub represents the LLM fine-tuned on the subset that filtered from the training set by the selection methods while the base LLM often involves 2 types: i)fine-tuned on the full training set and ii) fine-tuned on the same-scale subset filtered by the regular selections (e.g. random sampling and instruction length). In the formular 10, Num(win) represents the number of wining cases, Num(lose) represents the number of losing cases and Num(all) represents the number of all cases in the testing benchmarks.
The outputs of LLM-sub and the base LLM are rated by the judge on a scale from 1 to 10, and usually GPT4 is employed as the judge. To address the positional bias of the judge, Li et al. (2023a) send the outputs of these two LLMs to the judge twice with different orders. According to Li et al. (2023a), the wining cases refer to LLM-sub outperforming the base LLM in both times or winning in one and tying in the other. The losing cases refer to LLM-sub lagging behind the base LLM in both times or tying in one and losing in the other. The wining rates of different selection methods on the testing benchmarks are summarized in Table 1.
4.2 Inner Comparison
4.3 External Comparison
@@
4.4 Result Analysis
The proposed selection methods outperform the regular selections, which proves the significance of the data selection in instruction tuning. As shown in Tabel 1 and 2, TAGLM-13b-v1.0 and IFD outperform the regular selections on MT-benc based on instruction length and random sampling when tuning llama-13b on the mixture dataset, and Alpagsaus and IFD outperform the random sampling on total WS when tuning llama-7b on alpaca.
@@ 4.4 Result Analysis
5 Conclusions and Open Challenges
This paper provides a comprehensive overview of methods and challenges on data selection of instruction tuning, emphasizing the crucial role of high-quality data on the fine-tuning stage. We showcase some existing datasets along with their corresponding construction methods. These datasets suffer from issues such as imbalanced data distribution and inconsistent data quality. Based on this foundation, we present the existing four types of data selection methods. Methods relying solely on metric sets could effectively assess the quality of individual data, but lack consideration of complex dataset characteristics. Trainable large model methods that design scoring formulas could choose data suited to the LLM itself, and Methods using external powerful LLMs like GPT-4 for scoring have exceptional data selection capabilities. Furthermore, methods utilizing small models usually designed with multiple modules, considering various aspects. Finally we introduce how to evaluate the data selection methods, including wining rate, inner comparison and external comparison. Despite existing methods achieving commendable performance, several challenges persist.
There is a notable lack of uniform evaluation standards. In Section 4, we have presented multiple evaluation methods and various benchmarks concerning LLMs. Various data selection methods usually choose various evaluation standard, which has caused considerable difficulty in determining which method has more advantages. Future research could be devoted to establishing a reasonable, comprehensive, and automatic evaluation method for selection methods, thereby unifying the evaluation process for methods.
Handling large volumes of data often proves to be inefficient and exhibit a heavy reliance on powerful LLMs. When the scale of the instruction dataset to be filtered is too large, existing methods generally take too long to process, especially when using LLMs to extract instruction features or perform related metric calculations. Processing hundreds of thousands instruction data with LLMs can be very time-consuming.Furthermore, when using powerful LLMs like GPT4 API to process hundreds of thousands of instructions, the cost is very high. Future work should explore the use of smaller models and strive to match the selection capabilities of LLMs.
Existing models and methods for evaluating data quality primarily focus on English and general domains, lacking models for other languages and selection methods in specific domains.Future research should explore instruction quality assessment models for other languages, and investigate the design of task-specific selection methods. This aims to enhance the performance of methods across various scales, domains, and languages of instruction sets.
The challenges provide promising directions for further exploration, potentially leading to breakthroughs.In conclusion, it is our hope that these analytical efforts and the provided overview can further contribute to enhancing the abilities of LLMs in following human instructions.

### v2

@@ 4.2. Inner Comparison
4.3. External Comparison
@@
External comparison helps reveal the generalization ability of data selection methods across different base models and architectures. For example, as shown in Table 4, IFD(Li et al., 2024c) demonstrates robust performance when transferring from Llama-7B to Llama2-7B, indicating that the selected data subsets contain general instruction information beneficial to multiple models. However, this evaluation also exposes some limitations: certain methods that perform well in inner comparison show diminished advantages in external comparison, possibly due to overfitting to the training distribution of specific models. For instance, although Alpagsaus (Chen et al., 2024) performs excellently in inner comparison, its performance is unstable when compared against multiple external LLMs.
4.4. Result Analysis
The proposed selection methods outperform the regular selections, which proves the significance of the data selection in instruction tuning. As shown in Tabel 2 and 3, TAGLM-13B-v1.0 and IFD outperform the regular selections on MT-benc based on instruction length and random sampling when tuning Llama-13B on the mixture dataset, and Alpagsaus and IFD outperform the random sampling on total WS when tuning Llama-7B on alpaca.
@@ 4.4. Result Analysis
The above findings indicate that no single method holds an absolute advantage across all tasks. Instead, hybrid approaches—such as using metrics for preliminary filtering followed by detailed evaluation by large language models (LLMs)—may represent a more promising direction. Future evaluations should further take into account the characteristics of downstream tasks and the types of instructions, rather than relying solely on coarse-grained metrics like overall accuracy or win rate.
5. Discussion and Open Challenges
This paper provides a comprehensive overview of methods and challenges in instruction data selection for LLM tuning, emphasizing the crucial role of high-quality data during the fine-tuning stage. We review existing instruction datasets and their construction approaches, highlighting common issues such as imbalanced data distribution and inconsistent quality. On this basis, we categorize four representative types of data selection methods. Indicator-based methods can effectively assess sample-level quality but often overlook dataset-level characteristics. Trainable LLM-based methods adapt to model-specific preferences through scoring functions. Methods based on powerful external LLMs (e.g., GPT-4) offer superior general scoring capacity. Lastly, small model-based methods often adopt modular pipelines to approximate high-cost selections efficiently.
We also introduce three evaluation paradigms: win rate (against reference data), inner comparison (vs. full/small subsets with the same model), and external comparison (vs. other LLMs). Although existing methods show promising results, several challenges remain:
Lack of unified evaluation standards. As detailed in Section 4, different methods adopt different evaluation schemes and benchmarks, making it difficult to compare across techniques fairly. Future research should aim to develop standardized, automatic evaluation protocols that comprehensively assess selection quality under consistent conditions.
High computational cost and dependence on large models. When filtering hundreds of thousands of instructions, methods using large LLMs or APIs (e.g., GPT-4) can be prohibitively expensive and slow. There is a pressing need for efficient selection approaches that reduce reliance on external models, for example, by leveraging distilled or lightweight selectors that mimic the behavior of larger models.
Limited coverage across domains and languages. Most current selection frameworks are designed for English and general-purpose tasks. Extending them to other languages (e.g., Chinese, Arabic) and specialized domains (e.g., biomedical, legal) remains an open challenge.
To concretize the future directions above, we propose several specific and actionable research paths:
First, addressing the evaluation inconsistency, future work should build a standardized benchmark suite for instruction selection. This includes diverse downstream tasks (reasoning, summarization, code generation), a variety of model architectures (e.g., LLaMA, Mistral), and multilingual datasets. In addition, we advocate for task-conditioned evaluation, where selection quality is judged by its impact on real-world applications rather than just coarse-grained win rates.
Second, to improve efficiency and scalability, a promising strategy is to develop distillable selector models—compact models trained to emulate powerful LLMs’ selection behavior. Another approach is retrieval-augmented pruning, in which fast heuristics narrow down candidate data, followed by LLM-based fine filtering only on top-ranked items, significantly reducing total cost.
Third, to enhance domain and language adaptability, future methods can integrate multilingual encoders (e.g., XLM-R) and incorporate domain-specific priors, such as metadata or instruction patterns. Modular design would allow swapping components to suit medical, legal, or programming domains, supporting cross-lingual and cross-domain generalization.
Lastly, we strongly encourage the open-sourcing of reproducible selection pipelines, including code, pretrained selector models, scoring results, and evaluation benchmarks. This will ensure fair comparisons and promote rapid iteration across the community.
In summary, instruction data selection is a critical but underexplored component in the instruction tuning pipeline. We hope this survey and analysis serve as a solid foundation to guide further research and practical improvements in making LLMs more instruction-following and efficient.
Acknowledgements.
This work is supported by National Key R&D Proram of China (No.2022ZD0160602 and No.2022YFF0902703), the National Natural Science Foundation of China (No.62472121), and the Special Funding Program of Shandong Taishan Scholars Project.